Use this information to answer the question below:

  • The "spot rate" (current exchange rate) is Z2.01/$
  • The one-year forward rate is Z2.03/$
  • The annual risk-free borrowing/investing rate in the U.S. is 2.92%
  • The annual risk-free borrowing/investing rate in Country Z is 5.91%

In one round of arbitrage trades, how much money can you make by borrowing in one market and simultaneously investing in the other? Assume that you are starting out with $1,000,000 and that there are no transaction costs. Round your answer to the nearest dollar. The correct answer could be zero, but it's highly unlikely. Don't use dollar signs or commas in your answer.
Answer #1

Step 1). Borrow $1,000,000. After 1 year, you need to pay back 1,000,000*(1+ U.S. borrowing rate) =

1,000,000*(1+2.92%) = 1,029,200

Step 2). Convert $1,000,000 to Z at the spot rate of Z2.01/$

Amount (in Z) = 1,000,000*2.01 = 2,010,000

Step 3). Lend Z2,010,000 at the annual investing rate of 5.91% so after 1 year, you have

2,010,000*(1+5.91%) = Z2,128,791

Step 4). Convert Z to $ at the forward rate of Z2.03/$ to get 2,128,791/2.03 = $1,048,665.50

Step 5). Net profit (in $) = $1,048,665.50 - $1,029,200 = $19,465.52